OpenStudy (anonymous):

find the slope between these two points: (3,1) (5,4) (leave the answer as a fraction)

OpenStudy (anonymous):

\[slope=(y2-y1)/(x2-x1)\]

OpenStudy (amistre64):

subtract the points and stack y/x and reduce

OpenStudy (amistre64):

(5,4) -(3,1) ------ (2,3) ; slope = y/x = 3/2

jimthompson5910 (jim_thompson5910):

m=(y2-y1)/(x2-x1) m=(4-1)/(5-3) m=(3)/(5-3) m=3/2

OpenStudy (amistre64):

i subtract vertically so that I can keep track of it all :)

OpenStudy (amistre64):

dunno, i never go in there much

OpenStudy (anonymous):

do u always subtract?

jimthompson5910 (jim_thompson5910):

yes, you subtract the corresponding coordinates

OpenStudy (anonymous):

thanks